### v1

# Efficient Sampling in Disease Surveillance through Subpopulations: Sampling Canaries in the Coal Mine

Abstract
We consider disease outbreak detection settings where the population under study consists of various subpopulations available for stratified surveillance. These subpopulations can for example be based on age cohorts, but may also correspond to other subgroups of the population under study such as international travellers. Rather than sampling uniformly over the entire population, one may elevate the effectiveness of the detection methodology by optimally choosing a subpopulation for sampling. We show (under some assumptions) the relative sampling efficiency between two subpopulations is inversely proportional to the ratio of their respective baseline disease risks. This leads to a considerable potential increase in sampling efficiency when sampling from the subpopulation with higher baseline disease risk, if the two subpopulation baseline risks differ strongly. Our mathematical results require a careful treatment of the power curves of exact binomial tests as a function of their sample size, which are erratic and non-monotonic due to the discreteness of the underlying distribution. Subpopulations with comparatively high baseline disease risk are typically in greater contact with health professionals, and thus when sampled for surveillance purposes this is typically motivated merely through a convenience argument. With this study, we aim to elevate the status of such “convenience surveillance” to optimal subpopulation surveillance.
Keywords: Sampling efficiency; Outbreak Detection; Binomial testing; Traveller surveillance. MSC2020 subject classifications: 62D05, 62F03, 62P10
1 Context and Motivation
Disease outbreak detection is a critical component in public health management (Lombard and Buckeridge, 2007) and as a result a substantial body of research has been dedicated to the development and implementation of various outbreak detection methodologies based on public health data (see e.g. Dato et al. (2004); Unkel et al. (2012)). However, the effectiveness of such methods hinges crucially on the data sampling strategy, which has received far less attention. This lack of scrutiny leads to inefficient sampling practices and unwarranted pessimistic attitude towards common convenience sampling approaches. Motivated by this pessimism, we formally discuss efficient sampling strategies in the context of disease outbreak detection.
For the practitioner the headline of our discussion is that if subpopulations are available for stratified surveillance, then for the purposes of outbreak detection it can be much more efficient (i.e. requiring fewer samples for equal outbreak detection performance) to sample from a subpopulation rather than using a representative sample from the entire population. Specifically, sampling from subpopulations in which subjects have, under non-outbreak settings, the highest risk of contracting the disease under study may be considerably more efficient if their risk is relatively large in comparison to the average population risk. Such high-risk subpopulations potentially represent proverbial “canaries in the coalmine”, exhibiting evidence for disease outbreaks more clearly than the full population.
Our discussion spans active surveillance settings (Murray and Cohen, 2017). We start with a static setting where we sample subjects from a population, uncovering if the subject is currently carrying a disease under study, in order to infer if the disease prevalence in the population exceeds a known non-outbreak baseline prevalence. We then show how our insights from the static setting extend to a monitoring setting, where a sample of subjects is monitored over time, and data arrives sequentially in the form of daily new cases within this sample. In this setting the goal is to raise an alarm when the average rate of subjects getting diseased in the population exceeds a known baseline rate.
We consider outbreak detection settings where the population under study consists of multiple subpopulations (which may overlap) which we index by $j=1,\dots,s$, of size $m_{j}$, each available for stratified surveillance, and each with a subpopulation-specific baseline disease risk. Outbreaks are assumed to manifest as elevated number of disease cases across all subpopulations and proportional to the baseline disease risk. We revisit and generalize this assumption in more detail in the subsequent sections.
We characterize the sampling efficiency between samples from distinct subpopulations comprehensively. For the static surveillance setting a suitable outbreak detection methodology uses the exact binomial test for proportions. Our main contribution and central to our discussion lies in a novel non-asymptotic analytic quantification of the relative sampling efficiency between such exact binomial tests. The ensuing analysis is nontrivial due to the non-monotonic and erratic nature of the binomial test power curve as a function of the sample size (Chernick and Liu, 2012). We avoid normal approximations, as these have poor accuracy our regime of interest. The obtained relative sampling efficiency implies (under some assumptions) it is more efficient to sample from subpopulations with a high baseline prevalence. We show how this efficiency result can be extended to a monitoring setting under a common distributional approximation. Finally, an analysis of COVID-19 data further expands on the discussed model in a monitoring setting, corroborating applicability and potential gains of such sampling efficiency in practical settings.
Apart from potentially being more efficient, subpopulations with a high baseline risk of contracting a disease are typically more readily available for sampling as well, as individuals at elevated health risk are typically in greater contact with health professionals. For this reason, our insights are practically relevant and applicable in various contexts, for example:
  • Travellers to tropical destinations. In the context of tropical disease surveillance, travellers to tropical destinations can (temporarily) be seen as a subpopulation (in our statistical model) of their destination country. Travellers potentially have an elevated baseline risk of disease, due to increased close contact with other individuals (Freedman and Leder, 2006) or low natural resistance to diseases endemic in their destination (Wilson, 2003). This implies the baseline risk of contracting tropical diseases may be larger in the traveller subpopulation than in the domestic population. Therefore the subpopulation of travellers is a prime example of a subpopulation that is potentially much more efficient to monitor for the purposes of outbreak detection when compared to the domestic population itself. This contradicts the common pessimistic sentiment that monitoring travellers (as opposed to monitoring the full population) is merely a convenient monitoring approach (Fukusumi et al., 2016; Taylor-Salmon et al., 2023; Hamer et al., 2020; Leder et al., 2017; Bao et al., 2022).
  • Monitoring of low-risk diseases. When monitoring endemic diseases with low individual risk (e.g. influenza or COVID-19) many individuals do not communicate infections to health care professionals (Lau et al., 2021; Wang et al., 2022), thereby obfuscating potential upward trends in case numbers. Instead of focusing resources on getting overall reporting rate up, it may be more efficient to focus efforts on elevating reporting rates within subpopulations for which baseline risk of contracting disease is known to be highest. We discuss a real-world COVID-19 example that underscores this in Section 4.
There are only a few works that formally discuss sampling strategy in the context of disease surveillance. In Ferguson et al. (2014), optimal sampling is discussed in zoonotic disease epidemics where the goal is to balance sampling between hosts and vectors. In Murato et al. (2020) spatial aspects of sampling within pig farms are discussed with the main goal of devising sampling strategies that are not labor intensive.
2 Static Population Sampling
Consider a static setting where an unknown number of subjects within a population is carrying a disease and the goal is to infer, through a sample from the population, if the proportion of diseased subjects is larger than some pre-defined maximum baseline proportion. Our goal is to identify which subpopulation would be most efficient for this purpose. In the context of this static setting, subpopulations with a higher disease risk will have a larger proportion of diseased subjects.
Denote with $x_{i,j}$ the disease status of subject $i$ from subpopulation $j$, with $x_{i,j}=1$ if subject $i$ carries the disease and zero otherwise. Denote the fraction of subjects carrying the disease within subpopulation $j$ under a non-outbreak scenario by $p_{j}$ (i.e. the baseline prevalence) which we assume is known. Denote the unknown fraction of diseased subjects by $q_{j}=m_{j}^{-1}\sum_{i=1}^{m_{j}}x_{i,j}$. Our aim is to test, for a subpopulation $j$, for the following pair of competing hypotheses:
$H_{0}^{(j)}:q_{j}\leq p_{j}\text{ vs. }H_{1}^{(j)}:q_{j}>p_{j}\ .$ (1)
We make no assumptions on the risk profiles of subjects $x_{i,j}$ at an individual level, but we assume these are unknown and cannot be leveraged for sampling purposes. Therefore for our testing problem in (1) one can do no better than sampling uniformly over the subpopulation. Subsequently, this heterogeneity of risk profiles plays no role in the distribution of the sampled observations, reducing the problem to a binomial testing scenario. Consider a sample of size $n$ of observations from subpopulation $j$ as $\{X_{1,j},\dots,X_{n,j}\}$ where $X_{i,j}\sim\text{Bernoulli}(q_{j})$. The quantity of interest is now, for each subpopulation, the minimum required sample size for the exact binomial test to have power under the alternative at least some prescribed $1-\beta$ conducted at significance level $\alpha$. Naturally, it is then most efficient to sample from the subpopulation with lowest sample size requirements.
At this point, one may be tempted to proceed with arguments under normal approximations. However, the present context implies we are interested in small values of $p_{j}$ and $q_{j}$ which make such approximations inaccurate even at sample sizes larger than commonly understood sufficient; see Brown et al. (2001, 2002); Chernick and Liu (2012) for further discussion. We therefore do not make these approximations here, and our technical arguments must carefully treat the underlying non-monotonic and erratic nature of the exact binomial test power curve as a function of the sample size (Chernick and Liu, 2012).
We first restrict ourselves to two subpopulations where the baseline prevalence is higher in subpopulation $1$, i.e. $p_{1}>p_{2}$. Let $\psi_{j,n}:\mathbb{R}^{n}\to\{0,1\}$ denote the exact binomial test at significance level $\alpha$ for the pair of hypotheses in (1) using a sample of size $n$ from subpopulation $j$. Denote the power of this test at significance level $\alpha$ under the alternative where $q_{j}>p_{j}$ by $\mathcal{P}_{\alpha}(\psi_{j,n})$. We have the following result pertaining relative sampling efficiency between tests conducted on the two subpopulations:
Theorem 1.
Let $n_{2}\in\mathbb{N}$, and $p_{2}\leq q_{2}\leq 1/2$ be such that
$\alpha\leq\tfrac{1}{2}-\sqrt{\frac{1}{2en_{2}p_{2}(1-p_{2})}}\ ,\quad\mathcal{P}_{\alpha}(\psi_{2,n_{2}})\geq\tfrac{1}{2}+\sqrt{\frac{2}{en_{2}q_{2}(1-q_{2})}}\ .$ (2)
If
$n_{1}=\frac{p_{2}}{p_{1}}\cdot n_{2}\ ,$ (3)
and
$\frac{q_{1}}{q_{2}}\geq\frac{p_{1}}{p_{2}}\ ,$ (4)
then
$\mathcal{P}_{\alpha}(\psi_{1,\lfloor n_{1}\rfloor})>\mathcal{P}_{\alpha}(\psi_{2,{n_{2}}})-\sqrt{\frac{q_{1}}{2e\lfloor n_{1}\rfloor(1-q_{1})}}\mathbbm{1}\left\{n_{1}\text{ is not integer }\right\}\ .$ (5)
The proof is deferred to Section A.1. Theorem 1 contrasts the sample size requirements for the tests on the two subpopulations. Specifically, given a sample size for the test conducted on subpopulation 2 with a low baseline prevalence ($n_{2}$), it quantifies how small the sample size from subpopulation 1 with a high baseline prevalence can be chosen (i.e. (3)), for the two tests to retain nearly (see Remark 1) identical power. In other words, it delineates assumptions needed such that for nearly identical testing power, the required sample size when sampling from the subpopulation with high baseline prevalence is a fraction ($p_{2}/p_{1}$) of the sample size required when sampling from the subpopulation with low baseline prevalence. This implies sampling from the subpopulation with high baseline prevalence is considerably more efficient if the ratio of baseline prevalences $p_{2}/p_{1}$ is small.
Note that (4) holds if we assume that the fraction of diseased subjects $q_{j}$ are the product of the baseline prevalence $p_{j}$ and a multiplicative factor $\nu>1$ which is identical over all subpopulations:
$q_{j}=\nu p_{j}\ .$ (6)
In a setting with multiple subpopulations, Theorem 1 informs how the sampling efficiency between subpopulations can be pairwise compared. A natural question is if any efficiency can be gained by combining multiple subpopulations. Under the multiplicative model for the outbreak proportion (6) this is impossible: since the baseline prevalence of a combination of subpopulations cannot exceed the highest baseline prevalence of its components, Theorem 1 implies the combined subpopulation is less efficient that the component with the highest baseline prevalence.
Remark 1.
Ideally, negative term in (5) would not be present and the testing power of the two tests would instead be identical. Its inclusion arises from the erraticness and non-monotonicity of the power curves (as a function of the sample size) of the exact binomial test, rooted in its discreteness. This affects practice and analysis even at large sample sizes; see Brown et al. (2001, 2002); Chernick and Liu (2012) for further discussion.
Remark 2.
It may seem one could reinterpret the statement of Theorem 1 to concern the uniformly most powerful test, by recognizing the exact binomial test is equivalent to a likelihood ratio test, and subsequently applying the Neyman-Pearson lemma. This interpretation is incorrect, however, since this likelihood ratio test has exact type-I error only at a discrete set of significance levels. The uniformly most powerful test in this context is a randomized test, which by nature are typically undesirable in practice and thus excluded from our discussion.
3 Monitoring Cases in a Sample
We extend the static setting of the previous section to a monitoring setting to show how the previously obtained results may be applied and generalized. As before, our aim is to describe the relative sampling efficiency between samples from distinct subpopulations. Unlike the static setting where each subject $i$ in subpopulation $j$ has a fixed diseased status $x_{i,j}$, we now assume that each subject has a recurrent (potentially subject-specific) probability of getting diseased. For the purposes of exposition, we assume a daily probability of getting diseased, but any context-specific appropriate monitoring frequency can be substituted. We denote this probability by $\omega_{i,j}$. In the context of this monitoring setting, subpopulations with a higher disease risk will have subjects with larger daily probabilities of getting diseased.
We observe the sequence of all daily newly diseased subjects within a sample of size $n_{j}$ from subpopulation $j$, which we assume is sampled uniformly at random prior to the monitoring period, and remains fixed throughout the monitoring period. Denote the set of sampled subject indexes by $I_{j}$ and the sequence of daily cases from this sample $I_{j}$ by $D_{j,1},D_{j,2},\dots|I_{j}$. If we assume no dependencies between the sampled subjects, and assume subjects may be diseased multiple times over the time horizon, then for all $t\in\mathbb{N}$ the daily observed cases can be modelled as a sum of Bernoulli variables, i.e. $D_{j,t}|I_{j}=\sum_{i\in I_{j}}B_{i,j}$ with $B_{i,j}\sim\text{Bernoulli}(\omega_{i,j})$. In this setting, one wishes to use the sequence of observed cases $D_{j,1},D_{j,2},\dots|I_{j}$ to infer if the average rate of subjects getting diseased within subpopulation $j$, which we denote by $\omega_{j}=m_{j}^{-1}\sum_{i=1}^{m_{j}}\omega_{i,j}$, is larger than some pre-specified non-outbreak rate $\lambda_{j}$. In other words, we wish to test if
$H_{0}^{(j)}:\omega_{j}=\lambda_{j}\text{ vs. }H_{1}^{(j)}:\omega_{j}>\lambda_{j}\ .$ (7)
With knowledge of the sample size $n_{j}$, testing the above hypothesis through a sequential application of the exact binomial test may be appropriate. Such an approach then directly inherits the sampling efficiency given by Theorem 1, implying the same sampling efficiency conclusions in this monitoring setting.
However, in such settings one may not have knowledge of $n_{j}$ and $\lambda_{j}$ individually, but only of their product $n_{j}\lambda_{j}$ (i.e. the mean daily cases of the sample of size $n_{j}$ under non-outbreak circumstances). Methodology is then typically based on Poisson distributions; for example Poisson CUSUM control charts (Brook and Evans, 1972; Rogerson and Yamanda, 2004; Jiang et al., 2013). This approach can be motivated through the following approximations:
$D_{j,t}|I_{j}=\textstyle\sum_{i\in I_{j}}B_{i,j}\approx P\Big{(}\textstyle\sum_{i\in I_{j}}\omega_{i,j}\Big{)}\approx P\left(n_{j}\omega_{j}\right)\ ,$ (8)
where $P(\mu)$ denotes a Poisson random variable with mean $\mu$. In (8), the first approximation is due to Le Cam (1960). In the second approximation, the random rate (induced by the sample $I_{j}$) is replaced by its expectation. Note that under this approximation the involved distribution only depends on the product $n_{j}\lambda_{j}$ under the null. The CUSUM statistic can be defined recursively as
$C_{t}=\text{max}\left(0,C_{t-1}+D_{j,t}-k\right)\ ,$
where $k>0$ is a reference value typically chosen such that $C_{t}$ is a scaled version of Wald’s sequential probability ratio test (Lorden, 1971; Lucas, 1985). The null hypothesis in (7) is rejected (and an outbreak alarm raised) whenever $C_{t}$ exceeds a prescribed control limit $h$. The run length of a control chart is defined as the time until the chart signals an alarm. Now, one aims to calibrate $h$ such that the average run length is suitably large under the null, while being as small as possible under the alternative.
Under the approximations (8), one can directly relate the distributions of the sequential samples to obtain results analogous to those of Theorem 1 for the Poisson CUSUM control chart. As in the static setting, consider two subpopulations where the baseline average disease rate is larger in subpopulation 1, i.e. $\lambda_{1}>\lambda_{2}$. Consider sampling efficiency and outbreak rate assumptions analogous to the static setting (3) and (4) respectively as:
$n_{1}=\Big{\lfloor}\frac{\lambda_{2}}{\lambda_{1}}\cdot n_{2}\Big{\rfloor}\ ,\quad\frac{\omega_{1}}{\omega_{2}}\geq\frac{\lambda_{1}}{\lambda_{2}}\ .$ (9)
[…]
Thus under the reduced sample choice (9), observations under the null from subpopulation 1 are stochastically dominated by observations from subpopulation 2. This stochastic domination implies that alarm thresholds $h$ which bound the average run length of the CUSUM statistic under the null for observations from subpopulation 2, also conservatively bound that same quantity for the CUSUM statistic from the reduced sample from subpopulation 1 (see Heidema et al. (2024) for a detailed proof). Under the alternative hypothesis where $\omega_{j}>\lambda_{j}$ we instead have
[…]
where first stochastic domination is due to (9) and the second due to the bound $\lfloor x\rfloor/x\geq 1-1/x$. The final approximation in (11) is due to rounding in (9) and holds exactly if the non-rounded quantity in (9) is an integer. The above implies under the alternative in (7), the observations from the small sample from subpopulation 1 (approximately) stochastically dominate the observations from the large sample from subpopulation 2. Since the alarm thresholds can be chosen identically by the arguments from (10) this implies (if the approximation would above would be exact) the average run length is smaller for the observations from subpopulation 1. A sharp exact characterization of how these approximations, as a result from rounding errors, propagate to the average run length is outside of the scope of this manuscript. Such an analysis is challenging due to the sequential dependencies of the CUSUM statistic. Related literature merely analyses this average run lengths numerically, typically through implicit formulations following Markov chain constructions (Lucas, 1985; Hawkins and Olwell, 1998), and remains an interesting avenue for future work.
4 COVID-19 Case Study
We further highlight the usefulness of the above insights through a case study of daily COVID-19 infections. As in the previous section, we consider monitoring daily infections in an ongoing fashion. We further expand on the setting and methodology, and consider an approach that can account for potential overdispersion of the observed counts (i.e. where the approximation in (8) is inaccurate). A suitable methodology is the negative binomial Generalized Likelihood Ratio (GLR) control chart (Höhle and Paul, 2008). This extension highlights how the insights from preceding sections continue to hold when more complex monitoring strategies are employed.
We consider daily COVID-19 infections in the Netherlands between the first of June, 2020 and the end of August 2020111Retrieved from https://data.rivm.nl/meta/srv/dut/catalog.search#/metadata/2c4357c8-76e4-4662-9574-1deb8a73f724. At the start of this period, free PCR testing for anyone with mild symptoms was made available and the number of cases was low and stable [^2] See https://www.rivm.nl/en/novel-coronavirus-covid-19/archive-covid-19-updates. It was relevant to monitor the case numbers closely for evidence of another infection wave potentially warranting interventions.
We consider the period between June 1st and July 6th as a baseline period, which we use to estimate the relevant baseline rate parameters which we will subsequently assume known for monitoring starting from the 7th of July. For the purposes of exemplifying the insights of the preceding sections, we contrast the effectiveness of monitoring based on two distinct subpopulations. As subpopulations, we consider two age groups; age 20-29 and age 40-49.
In this case study we do not have control over the sample size. Instead, since no age group was specifically targeted for testing, and both age groups are demographically of the same size [^3] See https://www.cbs.nl/en-GB/visualisaties/dashboard-bevolking/bevolkingspiramide, we assume the underlying samples are of the same size. The average daily infections during the baseline period are 19.5 and 15.7 for age group 20-29 and 40-49 respectively. Our insights on the sampling efficiency then imply (under assumptions akin to (9)) tests using age group 20-29, i.e. the subpopulation with the higher baseline rate, are more powerful and thus should detect outbreaks earlier. The case numbers are plotted [^4] Numerical studies are conducted using R (R Core Team, 2021) and the surveillance package (Salmon et al., 2016). in Figure 1.
Figure 1: Left: stratified daily COVID-19 case numbers. Right: GLR statistic. Horizontal lines indicate the alarm thresholds. Vertical lines indicate the time of first alarm.
The GLR monitoring statistic and suitable alarm thresholds are plotted in Figure 1. The statistic based on the subpopulation with high baseline rate (age 20-29) raises an alarm earlier than the statistic based on the subpopulation with low baseline rate (age 40-49), corroborating insights from the previous sections. The subpopulation with high baseline rate raises an alarm earlier at other values of the threshold as well; the plotted threshold simply serves as an example. In Figure 1 the alarm threshold is chosen such that the in-control average run length is 370 days, matching the Shewart control chart with default settings (Hawkins and Olwell, 1998).
In the present context authorities naturally do not exclusively obtain data from a single age group. Our conclusion above does not imply one should exclusively analyze data from subpopulations with the highest baseline rate; rather it implies that, if data is collected for the purposes of outbreak detection, collection efforts should be targeted at the subpopulation where one knows the baseline rate is largest.
Acknowledgments
I would like to thank R.C. Gündlach, S. Heidema, E.R. van den Heuvel, R. Huits, and R.A.J. Post for useful discussions. This project was funded through a cooperative agreement between the US Centers for Disease Control and Prevention and the International Society of Travel Medicine (Federal Award Number: 1U01CK000632-01-00).
Appendix A Appendix
A.1 Proof of Theorem 1
Proof.
The exact binomial test under study is given by:
$\psi_{j,n_{j}}=\mathbbm{1}\left\{\sum_{i=1}^{n_{j}}X_{i,j}>F_{n_{j},p_{j}}^{-1}(1-\alpha)\right\}\ ,$
where $F_{n,p}(x)$ denotes the cumulative distribution function (CDF) of the binomial distribution with count parameter $n$ and probability parameter $p$, and likewise $F_{n,p}^{-1}(u)$ denoting its quantile function. We must now suitably lower bound its power
$\mathcal{P}_{\alpha}\left(\psi_{1,\lfloor n_{1}\rfloor}\right)=1-F_{\lfloor n_{1}\rfloor,q_{1}}\left(F^{-1}_{\lfloor n_{1}\rfloor,p_{1}}(1-\alpha)\right)\ ,$
which can be done through upper bounding both $F_{\lfloor n_{1}\rfloor,q_{1}}(x)$ and $F^{-1}_{\lfloor n_{1}\rfloor,p_{1}}(1-\alpha)$. The proof hinges on bounds and monotonicity results of the binomial cumulative distribution function, where one must carefully account for the discreteness of the distributions involved. Before proceeding with the proof, we first state the technical results needed. A standard result is that
$F_{n,u}(x)\leq F_{n,v}(x)\ ,$ (12)
for all $x\in\mathbb{R}$ when $u\geq v$ through a stochastic domination argument. A crucial and more technical result is the following lemma, which is a minor adaptation from the results derived in Anderson and Samuels (1967):
Lemma 1 (adapted from Anderson and Samuels (1967)).
Let $n/a\in\mathbb{N}$ and $a\geq 1$. Then:
$F_{n,p}(x)$ $<F_{\frac{n}{a},ap}(x)\quad\forall x\geq np\ ,$
$F_{n,p}(x)$ $>F_{\frac{n}{a},ap}(x)\quad\forall x\leq np-1\ .$
The proof is provided in Section A.2. We also require a number of other technical results, which we collect in the following lemma and prove in Section A.3:
Lemma 2.
  1. (i)
    For all $x\in\mathbb{R}$:
    $F_{n,p}(x)-F_{(n+1),p}(x)\leq\sqrt{\frac{p}{2en(1-p)}}\ .$
  2. (ii)
    If $n/a\in\mathbb{N}$ and $a\geq 1$, then:
    $F_{n,p}^{-1}(u)\geq F_{\frac{n}{a},ap}^{-1}(u)\quad\forall u\geq\tfrac{1}{2}+\sqrt{\tfrac{1}{2enp(1-p)}}\ .$
  3. (iii)
    For all $n$ and $p$ we have
    $F^{-1}_{n,p}\left(\tfrac{1}{2}-\sqrt{\tfrac{2}{enp(1-p)}}\right)$ $\leq np-1\ ,$
    $F^{-1}_{n,p}\left(\tfrac{1}{2}+\sqrt{\tfrac{1}{2enp(1-p)}}\right)$ $\geq np\ .$
We start by bounding $F_{\lfloor n_{1}\rfloor,q_{1}}(x)$. First note if $x\leq n_{2}q_{2}-1$ we have
$F_{\lceil n_{1}\rceil,q_{1}}(x)=F_{\frac{n_{2}}{n_{2}/\lceil n_{1}\rceil},q_{1}}(x)\leq F_{\frac{n_{2}}{n_{2}/\lceil n_{1}\rceil},\frac{q_{2}n_{2}}{\lceil n_{1}\rceil}}(x)<F_{n_{2},q_{2}}(x)\ ,$ (13)
where the first inequality follows from (12) since assumptions (3) and (4) ensure
$\lceil n_{1}\rceil\geq n_{1}=\frac{p_{2}n_{2}}{p_{1}}\geq\frac{q_{2}n_{2}}{q_{1}}\Longleftrightarrow q_{1}\geq\frac{q_{2}n_{2}}{\lceil n_{1}\rceil}\ ,$
and the second inequality follows from application of the second result of Lemma 1 with $a=n_{2}/\lceil n_{1}\rceil\geq 1$. We can then bound our quantity of interest by observing that for $x\leq n_{2}q_{2}-1$:
$F_{\lfloor n_{1}\rfloor,q_{1}}(x)$ $=F_{\lceil n_{1}\rceil,q_{1}}(x)+\Bigg{(}F_{\lfloor n_{1}\rfloor,q_{1}}(x)-F_{\lfloor n_{1}\rfloor+1,q_{1}}(x)\Bigg{)}\mathbbm{1}\left\{\lfloor n_{1}\rfloor\neq\lceil n_{1}\rceil\right\}$
$<F_{n_{2},q_{2}}(x)+\sqrt{\frac{q_{1}}{2e\lfloor n_{1}\rfloor(1-q_{1})}}\mathbbm{1}\left\{\lfloor n_{1}\rfloor\neq\lceil n_{1}\rceil\right\}\ ,$ (14)
where we used (13) and result (i) from Lemma 2. Next, we bound the quantile $F_{\lfloor n_{1}\rfloor,p_{1}}^{-1}(1-\alpha)$ as
$F_{\lfloor n_{1}\rfloor,p_{1}}^{-1}(1-\alpha)=F_{\frac{n_{2}}{n_{2}/\lfloor n_{1}\rfloor},p_{1}}^{-1}(1-\alpha)\leq F_{\frac{n_{2}}{n_{2}/\lfloor n_{1}\rfloor},\frac{p_{2}n_{2}}{\lfloor n_{1}\rfloor}}^{-1}(1-\alpha)\leq F_{n_{2},p_{2}}^{-1}(1-\alpha)\ ,$ (15)
where the first inequality follows from (12) since assumption (3) ensures
$\lfloor n_{1}\rfloor\leq n_{1}=\frac{p_{2}n_{2}}{p_{1}}\Longleftrightarrow p_{1}\leq\frac{p_{2}n_{2}}{\lfloor n_{1}\rfloor}\ ,$
and the second inequality follows from result (ii) from Lemma 2 with $a=n_{2}/\lfloor n_{1}\rfloor\geq 1$ since (2) implies $1-\alpha\geq\tfrac{1}{2}+1/\sqrt{2en_{2}p_{2}(1-p_{2})}$. As our bound (14) is only valid for $x\leq n_{2}q_{2}-1$ we must ensure the bound on our quantile (15) is small enough such that we can use both bounds in unison. Since the power of the test $\mathcal{P}_{\alpha}(\psi_{2,n_{2}})\geq\frac{1}{2}+\sqrt{2/(en_{2}q_{2}(1-q_{2}))}$ by (2) we can upper bound the quantile $F_{n_{2},p_{2}}^{-1}(1-\alpha)$ by
$\mathcal{P}_{\alpha}(\psi_{2,n_{2}})\geq\frac{1}{2}+\sqrt{\frac{2}{en_{2}q_{2}(1-q_{2})}}$
[…]
where in the last inequality we used (iii) from Lemma 2. Having ensured (16), we ultimately find
$\mathcal{P}_{\alpha}(\psi_{1,\lfloor n_{1}\rfloor})$ $=1-F_{\lfloor n_{1}\rfloor,q_{1}}\Big{(}F_{\lfloor n_{1}\rfloor,p_{1}}^{-1}(1-\alpha)\Big{)}$
$\geq 1-F_{\lfloor n_{1}\rfloor,q_{1}}\Big{(}F_{n_{2},p_{2}}^{-1}(1-\alpha)\Big{)}$
$>1-F_{n_{2},q_{2}}\Big{(}F_{n_{2},p_{2}}^{-1}(1-\alpha)\Big{)}-\sqrt{\frac{q_{1}}{2e\lfloor n_{1}\rfloor(1-q_{1})}}\mathbbm{1}\left\{\lfloor n_{1}\rfloor\neq\lceil n_{1}\rceil\right\}$
$=\mathcal{P}_{\alpha}(\psi_{2,n_{2}})-\sqrt{\frac{q_{1}}{2e\lfloor n_{1}\rfloor(1-q_{1})}}\mathbbm{1}\left\{\lfloor n_{1}\rfloor\neq\lceil n_{1}\rceil\right\}\ ,$
where the first inequality follows from (15) and the second inequality follows from (14) which holds due to the result in (16). ∎
A.2 Proof of Lemma 1
We adapt Theorem 2.1 from Anderson and Samuels (1967). Define $Y_{i}\sim\text{Bernoulli}(p_{i})$ for $i=1,\dots,n$. Denote the CDF of their sum by $H(k):=\mathbb{P}\left(\sum_{i=1}^{n}Y_{i}\leq k\right)$ and $\lambda=\sum_{i=1}^{n}p_{i}$. Now Hoeffding (1956) implies
$F_{n,\lambda/n}(k)$ $\geq H(k)\quad\forall k\leq\lambda-1\ ,$
$F_{n,\lambda/n}(k)$ $\leq H(k)\quad\forall k\geq\lambda\ ,$
where equality holds only if all $p_{i}=\lambda/n$. Particularize $\lambda=np$ with $p_{1},\dots,p_{n-n/a}=0$ and
$p_{n-n/a+1},\dots,p_{n}=\frac{np}{n/a}=ap\ .$
Under this particularization, $H(k)=F_{n/a,ap}(k)$, such that the results above imply the statement of the lemma.
A.3 Proof of Lemma 2
We start by proving statement (i). Denote $Y_{i}\sim\text{Bernoulli}(p)$, we have that for any $x\in\mathbb{R}$:
$F_{(n+1),p}(x)$ $=\mathbb{P}\left(\sum_{i=1}^{n+1}Y_{i}\leq x\right)$
$=\mathbb{P}\left(\sum_{i=1}^{n}Y_{i}\leq x-1\right)\mathbb{P}\left(Y_{n+1}=1\right)+\mathbb{P}\left(\sum_{i=1}^{n}Y_{i}\leq x\right)\mathbb{P}\left(Y_{n+1}=0\right)$
$=pF_{n,p}(x-1)+(1-p)F_{n,p}(x)$
$=p\Big{(}F_{n,p}(x)-f_{n,p}(\lfloor x\rfloor)\Big{)}+(1-p)F_{n,p}(x)$
$=-pf_{n,p}(\lfloor x\rfloor)+F_{n,p}(x)\ .$
Rearranging the quantity above and using that (Herzog, 1947)
$f_{n,p}(x)\leq\sqrt{\frac{1}{2enp(1-p)}}\ ,$ (17)
completes the proof. Next we prove statement (iii). We have that
$F_{n,p}(np-1)$ $=F_{n,p}(\lceil np\rceil)-\Big{(}F_{n,p}(\lceil np\rceil)-F_{n,p}(\lfloor np-1\rfloor)\Big{)}$
$\geq\tfrac{1}{2}-f_{n,p}(\lceil np\rceil)-f_{n,p}(\lfloor np\rfloor)$
$\geq\tfrac{1}{2}-2\sqrt{\frac{1}{2enp(1-p)}}\ ,$
where in the first inequality we used that the median of the binomial distribution is smaller than $\lceil np\rceil$, and in the second inequality we used (17). Rearranging the inequalities then implies the first result of statement (iii). Now, denote $Z_{n,p}$ a binomial random variable with count $n$ and probability $p$, such that analogously we obtain
$1-F_{n,p}(np)$ $=\mathbb{P}\left(Z_{n,p}>np\right)$
$=\mathbb{P}\left(Z_{n,p}\geq\lfloor np\rfloor+1\right)$
$=\mathbb{P}\left(Z_{n,p}\geq\lfloor np\rfloor\right)-f_{n,p}(\lfloor np\rfloor)$
$\geq\tfrac{1}{2}-\sqrt{\frac{1}{2enp(1-p)}}\ ,$
where we used that the median of the binomial distribution is larger than $\lfloor np\rfloor$ and (17). Rearranging the inequality implies the second result of statement (iii). Finally, we prove statement (ii). Recall
$F^{-1}_{n,p}(u)=\inf\{x\in\mathbb{R}:F_{n,p}(x)\geq u\}\ .$
Lemma 1 implies that
$\forall x\geq np:\quad F_{n,p}(x)<F_{n/a,ap}(x)$
[…]
Now assume that $u\geq 1/2+\sqrt{1/2enp(1-p)}$. Then setting $F_{n,p}(x)\geq u$ and solving for $x$ implies (through monotonicity of $F_{n,p}$ and statement (iii) that:
$F_{n,p}(x)\geq u\Rightarrow x\geq F^{-1}_{n,p}(u)\geq np\ .$
Therefore under the assumption $u\geq 1/2+\sqrt{1/2enp(1-p)}$ we have:
$\{x\geq np:F_{n,p}(x)\geq u\}=\{x\in\mathbb{R}:F_{n,p}(x)\}\ ,$
and likewise
$\{x\geq np:F_{n/a,ap}(x)\geq u\}=\{x\in\mathbb{R}:F_{n/a,ap}(x)\}\ ,$
Finally this implies, if $u\geq 1/2+\sqrt{1/2enp(1-p)}$ that
$F_{n,p}^{-1}(u)=\inf\{x\in\mathbb{R}:F_{n,p}(x)\}\geq\inf\{x\geq np:F_{n/a,ap}(x)\geq u\}=F_{n/a,ap}^{-1}(x)\ ,$
which completes the proof.